# Schrödinger Bridge Matching for Tree-Structured Costs and Entropic Wasserstein Barycentres

### v1

@@ Abstract
Recent advances in flow-based generative modelling have provided scalable methods for computing the Schrödinger Bridge (SB) between distributions, a dynamic form of entropy-regularised Optimal Transport (OT) for the quadratic cost. The successful Iterative Markovian Fitting (IMF) procedure solves the SB problem via sequential bridge-matching steps, presenting an elegant and practical approach with many favourable properties over the more traditional Iterative Proportional Fitting (IPF) procedure. Beyond the standard setting, optimal transport can be generalised to the multi-marginal case in which the objective is to minimise a cost defined over several marginal distributions. Of particular importance are costs defined over a tree structure, from which Wasserstein barycentres can be recovered as a special case. In this work, we extend the IMF procedure to solve for the tree-structured SB problem. Our resulting algorithm inherits the many advantages of IMF over IPF approaches in the tree-based setting. In the specific case of Wasserstein barycentres, our approach can be viewed as extending fixed-point approaches for barycentre computation to the case of flow-based entropic OT solvers.
Preprint. Correspondence to howard@stats.ox.ac.uk
@@ 1 Introduction
Transporting mass between two distributions is a ubiquitous problem with numerous applications in machine learning and beyond. Optimal Transport (OT) (Santambrogio, 2015; Peyré and Cuturi, 2019) provides a principled approach for such problems, by seeking to minimise the total cost of transportation according to a chosen cost function. Since the introduction of Sinkhorn’s algorithm (Cuturi, 2013) and more recent neural approaches (Makkuva et al., 2020), computational OT has seen great success across many domains such as biology (Schiebinger et al., 2019; Bunne et al., 2023), and extensively in machine learning (Genevay et al., 2018; Cuturi et al., 2019; Corenflos et al., 2021). Recently, ideas from the powerful flow-based approaches that have revolutionised generative modelling (Song et al., 2021; Peluchetti, 2022; Lipman et al., 2023; Liu et al., 2023; Albergo and Vanden-Eijnden, 2023) have been leveraged to solve the entropy-regularised dynamic OT problem, known as the Schrödinger Bridge (SB). Such approaches provide significant scalability advantages, enabling approximation of OT maps between high-dimensional continuous datasets such as image data. Early flow-based SB solvers were based on the classical Iterative Proportional Fitting (IPF) scheme (De Bortoli et al., 2021; Vargas et al., 2021; Chen et al., 2022), but such methods have since been superseded by those based on the Iterative Markovian Fitting (IMF) scheme (Shi et al., 2023; Peluchetti, 2023) due to its many superior properties.
@@
Beyond the standard OT problem between two marginals, the multi-marginal OT problem considers obtaining a joint coupling over multiple marginal distributions while minimising the total cost. Tree-structured costs are often considered (Haasler et al., 2021), as they frequently arise in practical applications while also allowing for improved scalability by leveraging the tree structure. Of particular significance are star-shaped trees, as these correspond to the prominent Wasserstein barycentre problem (Agueh and Carlier, 2011). Computing the Wasserstein barycentre is notoriously challenging (Altschuler and Boix-Adserà, 2022). Many successful approaches approximate the barycentre with a finite set of points, but these in-sample methods struggle to scale well as dimensionality increases. Alternative methods aim to provide continuous approximations to the barycentre (Li et al., 2020) using neural parameterisations, and often require multi-level optimisation procedures (Fan et al., 2021; Korotin et al., 2022). Of particular relevance to our work is the recent diffusion-based approach of Noble et al. (2023) which extends the IPF-based Diffusion Schrodinger Bridge (DSB) method of De Bortoli et al. (2021) to the tree-based SB setting, and is (to our knowledge) currently the only neural ODE/SDE approach for barycentre computation.
@@
In this work, we extend the IMF procedure to the tree-based SB problem (Haasler et al., 2021). Our TreeDSBM algorithm provides an IMF counterpart to the TreeDSB method from Noble et al. (2023), closing a clear gap in the existing literature (see Table 1) and translating the many benefits of IMF over IPF to the tree setting. For the specific case of Wasserstein barycentre computation, our algorithm can be viewed as extending the commonly used fixed-point approaches to the case of flow-based entropic OT solvers. We demonstrate significantly improved empirical performance over TreeDSB, and illustrate that flow-based barycentre solvers can offer competitive performance against alternative algorithms for continuous Wasserstein-2 barycentre computation.
@@
SB TreeSB (Haasler et al., 2021)
IPF DSB (De Bortoli et al., 2021) TreeDSB (Noble et al., 2023)
IMF DSBM (Shi et al., 2023; Peluchetti, 2023) TreeDSBM (Ours)
Figure 1: The two stages of the TreeIMF procedure. On this tree, the marginals at the leaf vertices $\mathcal{S}=\{1,3,5,6,7\}$ (blue) are fixed. The marginals at vertices $\mathcal{V}\backslash\mathcal{S}$ (red) are not fixed, and change during the procedure.
Notation
We denote by $\mathcal{T}$ a tree with an edge length function, $\mathcal{T}=(\mathcal{V},\mathcal{E},\ell)$, or its associated metric space via associating edges $e$ with the line segment of length $T^{e}=\ell(e)$ (see Section 2.3). We let $\mathcal{P}(\mathcal{C})=\mathcal{P}(C([0,T],{\mathbb{R}}^{d}))$ and $\mathcal{P}(\mathcal{C}_{\mathcal{T}}):=\mathcal{P}(C(\mathcal{T},{\mathbb{R}}^{d}))$ denote the space of path measures on the interval $[0,T]$ and $\mathcal{T}$ respectively. For the interval $[0,T]$, $\mathcal{M}$ denotes the set of Markov measures, and $\mathcal{R}({\mathbb{Q}})$ denotes the reciprocal class for a reference measure ${\mathbb{Q}}\in\mathcal{M}$ (see Definitions 2.2 and 2.1). For a path measure ${\mathbb{Q}}\in\mathcal{P}(\mathcal{C})$, ${\mathbb{Q}}_{t}$ denotes the marginal distribution at time $t$, and ${\mathbb{Q}}_{\cdot|0,T}$ denotes the diffusion bridge conditional on values at $0,T$. We use $\Pi=\Pi_{0,T}{\mathbb{Q}}_{\cdot|0,T}\in\mathcal{P}(\mathcal{C})$ to denote mixtures of bridges $\Pi(\cdot)=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}{\mathbb{Q}}_{\cdot|0,T}(\cdot|x_{0},x_{T})\Pi_{0,T}({\mathrm{d}}x_{0},{\mathrm{d}}x_{T})$. Similarly on the tree $\mathcal{T}$, $\mathcal{M}_{\mathcal{T}}$ denotes the set of Markov measures, and $\mathcal{R}_{\mathcal{S}}({\mathbb{Q}})$ denotes the reciprocal class for a reference measure ${\mathbb{Q}}\in\mathcal{M}_{\mathcal{T}}$ and observed vertices $\mathcal{S}\subset\mathcal{V}$ (see Definitions 3.1 and 3.2). For a path measure ${\mathbb{Q}}\in\mathcal{P}(\mathcal{C}_{\mathcal{T}})$, ${\mathbb{Q}}_{v}$ denotes the marginal at vertex $v\in\mathcal{V}$, ${\mathbb{Q}}_{\mathcal{V}}$ denotes the induced joint distribution over a set of vertices $\mathcal{V}$, and ${\mathbb{Q}}_{\cdot|\mathcal{S}}$ denotes the bridge process conditional on values at vertices $\mathcal{S}$. We denote mixtures of bridges as $\Pi=\Pi_{\mathcal{S}}{\mathbb{Q}}_{\cdot|\mathcal{S}}\in\mathcal{P}(\mathcal{C}_{\mathcal{T}})$, defined similarly to above. Finally, ${\mathrm{KL}}(\pi\|\tilde{\pi})$ denotes the Kullback-Leibler divergence between measures $\pi_{0},\pi_{1}$, and ${\mathrm{H}}(\pi)$ denotes the differential entropy.
@@ 2 Background
2.1 Schrödinger Bridges and Optimal Transport
We begin by reviewing the standard Schrödinger Bridge (SB) problem between two distributions, and its relation to Optimal Transport (OT). Given an initial and final measure $\mu_{0}$ and $\mu_{T}$ of a population, the SB problem aims to identify the most likely intermediate dynamics of the population under the assumption that the movement is driven by a stochastic reference process ${\mathbb{Q}}$. The resulting dynamic SB problem is defined as
@@ 2.1 Schrödinger Bridges and Optimal Transport
Under mild assumptions, the two problems are equivalent; the dynamic SB solution can be expressed as a mixture of bridges over the static SB coupling, $\mathbb{P}^{SB}=\Pi_{0,T}^{SB}{\mathbb{Q}}_{\cdot|0,T}$. For more details see Léonard (2013).
@@ 2.2 Iterative Markovian Fitting
In fact, the SB solution $\mathbb{P}^{SB}$ can be characterised as the unique path measure that is both Markov, and a mixture of bridges $\mathbb{P}^{SB}=\mathbb{P}_{0,T}^{SB}{\mathbb{Q}}_{\cdot|0,T}$, that has correct marginals $\mathbb{P}^{SB}_{0}=\mu_{0},\mathbb{P}^{SB}_{T}=\mu_{T}$ (Léonard, 2013). This property motivates the Iterative Markovian Fitting (IMF) procedure (Shi et al., 2023; Peluchetti, 2023), which solves for the SB solution by alternately projecting between Markovian processes and processes with the correct bridges. We refer to Shi et al. (2023) for full details of the IMF procedure, but recall here the basic presentation. We recall the following definitions.
@@
A path measure $\Pi\in\mathcal{P}(\mathcal{C})$ is in the reciprocal class $\mathcal{R}({\mathbb{Q}})$ of ${\mathbb{Q}}$ if it is a mixture of bridges of ${\mathbb{Q}}$ conditional on their values at the endpoints, $\Pi=\Pi_{0,T}{\mathbb{Q}}_{\cdot|0,T}$. For $\mathbb{P}\in\mathcal{P}(\mathcal{C})$, its reciprocal projection is defined to be the mixture of bridges according to its induced coupling, $\operatorname{proj}_{\mathcal{R}({\mathbb{Q}})}(\mathbb{P})=\mathbb{P}_{0,T}{\mathbb{Q}}_{\cdot|0,T}$.
@@
It can be shown that, under mild conditions, these definitions coincide with the following minimisation problems over path measures (Shi et al., 2023),
@@
The IMF procedure requires learning the vector field corresponding the Markovian projections $\operatorname{proj}_{\mathcal{M}}(\Pi)$. This is done by training a neural network $v_{\theta}$ with a bridge-matching loss objective (Peluchetti, 2022), for which the drift in (4) is the optimum,
@@
Traditionally, the standard way to solve the SB problem is via the Iterative Proportional Fitting (IPF) procedure (Fortet, 1940; Kullback, 1968; Rüschendorf, 1995), which minimises the KL divergence between subsequent iterations while alternating the endpoint measure that is fixed.
@@
This method was implemented using diffusion model-based approaches in De Bortoli et al. (2021), giving the Diffusion Schrödinger Bridge (DSB) method (see also Vargas et al. (2021), Chen et al. (2022)). While allowing for arbitrary reference measures ${\mathbb{Q}}$, this approach suffers from several shortfalls, such as only preserving both marginals at convergence (in comparison to IMF, which preserves both marginals at each iteration), as well as expensive trajectory caching, and ‘forgetting’ of the original reference measure (Fernandes et al., 2022). As a result, IPF approaches have largely been superseded by IMF approaches which avoid these issues.
@@ 2.3 Tree-Structured Schrödinger Bridge
In this work, we consider the tree-structured Schrodinger Bridge problem (Haasler et al., 2021). In particular, we extend the IMF procedure to the dynamic Tree SB setting of Noble et al. (2023).
@@
In order to explain the dynamic Tree SB problem, we first need to define stochastic processes on a tree. We will consider trees $\mathcal{T}=(\mathcal{V},\mathcal{E},\ell)$ with vertex set $\mathcal{V}$ and edge set $\mathcal{E}$, where $\ell:\mathcal{E}\rightarrow{\mathbb{R}}^{>0}$ is an edge-length function. One can extend the tree to a uniquely arcwise-connected metric space in the natural way; the arc connecting the two endpoints of an edge $e$ is identified with a line segment of corresponding length $T^{e}=\ell(e)$, and they are connected according to the graph structure. Via a slight abuse of notation we will also denote this metric space as $\mathcal{T}$. As such, we can define the space $C(\mathcal{T},{\mathbb{R}}^{d})$ of continuous paths from $\mathcal{T}$ into ${\mathbb{R}}^{d}$, and we let $\mathcal{P}(\mathcal{C}_{\mathcal{T}}):=\mathcal{P}(C(\mathcal{T},{\mathbb{R}}^{d}))$ denote the space of probability measures over such paths. We will present the methodology according to a directed tree $\mathcal{T}_{r}$ rooted at a vertex $r$, for which we can choose an ordered edge set $\mathcal{E}_{r}$ corresponding to a depth-first traversal (though we will often omit the dependence on $r$ in the notation).
@@
As in the standard case, using the chain rule for KL divergence gives (under mild conditions) equivalence of the two problems; the dynamic SB solution can be expressed as a mixture of bridges over the static SB coupling, $\mathbb{P}^{SB}=\Pi_{\mathcal{V}}^{SB}{\mathbb{Q}}_{\cdot|\mathcal{V}}$. In this work, we aim to solve for the $\mathrm{TreeSB}_{\text{dyn}}$ problem.
@@
This is precisely an entropy-regularised multi-marginal OT problem with a tree-structured quadratic cost function $c(x_{\mathcal{V}})=\sum_{(u,v)\in\mathcal{E}}\tfrac{1}{T^{(u,v)}}\lVert x_{u}-x_{v}\rVert_{2}^{2}$ and entropy-regularisation value $\varepsilon=2\sigma^{2}$ (see Appendix for an overview). Note that each weight is the reciprocal of the corresponding edge length.
Wasserstein Barycentres
Given $\ell$ measures $(\mu_{1},...,\mu_{\ell})$ and weights $(\lambda_{1},...,\lambda_{\ell})$ summing to 1, the Wasserstein-2 barycentre (Agueh and Carlier, 2011) is defined as $\nu^{*}=\operatorname*{arg\,min}_{\nu}\sum_{i}\lambda_{i}W_{2}^{2}(\mu_{i},\nu).$ There are many ways to define entropy-regularised Wasserstein barycentres (see the Appendix for an overview). The work of Chizat (2023) unifies several of these, combining $\varepsilon$-regularised OT costs with an entropic penalty term on the barycentre weighted by $\tau$. Noble et al. (2023) show that the following doubly-regularised barycentre problem

### v2

@@ Abstract
Recent advances in flow-based generative modelling have provided scalable methods for computing the Schrödinger Bridge (SB) between distributions, a dynamic form of entropy-regularised Optimal Transport (OT) for the quadratic cost. The successful Iterative Markovian Fitting (IMF) procedure solves the SB problem via sequential bridge-matching steps, presenting an elegant and practical approach with many favourable properties over the more traditional Iterative Proportional Fitting (IPF) procedure. Beyond the standard setting, optimal transport can be generalised to the multi-marginal case in which the objective is to minimise a cost defined over several marginal distributions. Of particular importance are costs defined over a tree structure, from which Wasserstein barycentres can be recovered as a special case. In this work, we extend the IMF procedure to solve for the tree-structured SB problem. Our resulting algorithm inherits the many advantages of IMF over IPF approaches in the tree-based setting. In the case of Wasserstein barycentres, our approach can be viewed as extending the widely used fixed-point approach to use flow-based entropic OT solvers, while requiring only simple bridge-matching steps at each iteration. Our code is available at https://github.com/samuel-howard/Tree_SB_Matching_Barycentres.
Corresponding author: howard@stats.ox.ac.uk
@@ 1 Introduction
Transporting mass between two distributions is a ubiquitous problem with numerous applications in machine learning and beyond. Optimal Transport (OT) [74, 69] provides a principled approach for such problems, by seeking to minimise the total cost of transportation according to a chosen cost function. Since the introduction of Sinkhorn’s algorithm [17] and more recent neural approaches [62], computational OT has seen great success across many domains such as biology [75, 10], and extensively in machine learning [35, 21, 16]. Recently, ideas from the powerful flow-based approaches that have revolutionised generative modelling [83, 67, 56, 57, 2] have been leveraged to solve the entropy-regularised dynamic OT problem, known as the Schrödinger Bridge (SB). Such approaches provide significant scalability advantages, enabling approximation of OT maps between high-dimensional continuous datasets such as image data. Early flow-based SB solvers were based on the classical Iterative Proportional Fitting (IPF) scheme [24, 89, 12], but such methods have since been superseded by those based on the Iterative Markovian Fitting (IMF) scheme [77, 68] due to its many superior properties.
@@
Beyond the standard OT problem between two marginals, multi-marginal OT aims to find a joint coupling over multiple marginals while minimising the total cost. Tree-structured costs are often considered [38], as they frequently arise in applications while also allowing for improved scalability by leveraging the tree structure. Of particular significance are star-shaped trees, as these correspond to the prominent Wasserstein barycentre problem [1]. The Wasserstein barycentre provides a natural notion of ‘average’ for probability distributions, and is widely studied due to its importance in applications, including in Bayesian learning [84], clustering [90], and representation learning [78] to name a few. Computing barycentres is notoriously challenging [3]. Many successful approaches approximate the solution with a finite set of points, but these in-sample methods struggle to scale well as dimensionality increases. Alternative methods aim to provide continuous approximations [54] using neural parameterisations, and often require multi-level optimisation procedures [28, 47]. Of particular relevance to our work is the diffusion-based approach of [64] which extends the IPF-based Diffusion Schrödinger Bridge (DSB) method of [24] to the tree-based SB setting, and is (to our knowledge) currently the only neural ODE/SDE approach for barycentre computation.
One of the most successful and elegant approaches for Wasserstein-2 barycentre computation is the iterative fixed-point approach, which involves iteratively updating a candidate barycentre $\nu$ by solving for the OT map from $\nu$ to each marginal $\mu_{i}$ and updating $\nu$ according to the induced coupling. This approach was popularised by the seminal work of [4] and has formed the basis of many algorithmic developments, including in machine learning [47]. It has strong performance, and has been observed to converge quickly in only a few iterations [55]. However, the procedure requires solving a complete OT problem to each marginal at each iteration, which is expensive as solving even a single OT problem can be challenging.
(a) Reciprocal process $\Pi$: The $Y_{\mathcal{S}}$ ($\times$) are sampled from the current coupling $\Pi_{\mathcal{S}}$ over $\mathcal{S}$. Conditional on the $Y_{\mathcal{S}}$, points ($\times$) at unknown marginals $\mathcal{V}\backslash\mathcal{S}$ are sampled as $Y_{\mathcal{S}^{c}|\mathcal{S}}\sim{\mathbb{Q}}_{\mathcal{S}^{c}|\mathcal{S}}$. Brownian bridges are drawn along the edges between the samples.
Figure 1: The two stages of TreeIMF. On this tree, the marginals at the leaf vertices $\mathcal{S}$ (blue) are fixed. The marginal at the vertex in $\mathcal{V}\backslash\mathcal{S}$ (red) is not fixed, and can change during the procedure.
@@
In this work, we extend the IMF procedure to the tree-based SB problem [38]. Our TreeDSBM algorithm provides an IMF counterpart to the TreeDSB method from [64], closing a clear gap in the existing literature (see Table 1) and translating the many benefits of IMF over IPF to the tree setting. For the specific case of Wasserstein barycentre computation, our algorithm can be viewed as extending the commonly used fixed-point approaches to the case of flow-based entropic OT solvers. In particular, we show that the iterations of the IMF scheme and the fixed-point barycentre solvers can be elegantly combined into a single iterative procedure, yielding a fixed-point-style algorithm that requires only inexpensive bridge-matching at each iteration. We demonstrate significantly improved empirical performance over TreeDSB, and show that flow-based barycentre solvers can offer competitive performance against existing algorithms for continuous Wasserstein-2 barycentre computation.
@@
SB TreeSB [38]
IPF DSB [24] TreeDSB [64]
IMF DSBM [77, 68] TreeDSBM (Ours)
@@ 2 Background
2.1 Schrödinger bridges, optimal transport, and Wasserstein barycentres
We begin by reviewing the standard Schrödinger Bridge (SB) problem between two distributions, and its relation to Optimal Transport (OT). For notation used in the paper, see Appendix A. Given an initial and final measure $\mu_{0}$ and $\mu_{T}$ of a population, the SB problem aims to identify the most likely intermediate dynamics of the population under the assumption that the movement is driven by a stochastic reference process $\mathbb{Q}$}. The resulting dynamic SB problem is defined as
@@ 2.1 Schrödinger bridges, optimal transport, and Wasserstein barycentres
Under mild assumptions, the two problems are equivalent; the dynamic SB solution can be expressed as a mixture of bridges over the static SB coupling, $\mathbb{P}^{SB}=\Pi_{0,T}^{SB}{\mathbb{Q}}_{\cdot|0,T}$ [60].
@@
Wasserstein barycentres
A key motivation for the tree-structured setting that we will consider is the important Wasserstein barycentre problem [1], which provides a natural notion of ‘average’ for probability distributions. Given $\ell$ measures $(\mu_{1},...,\mu_{\ell})$ and weights $(\lambda_{1},...,\lambda_{\ell})$ summing to 1, the Wasserstein-2 barycentre is defined as
where $W_{2}^{2}(\mu_{i},\nu)$ denotes the optimal transport cost between $\mu_{i}$ and $\nu$ for the quadratic ground-cost.
The Wasserstein barycentre problem is notoriously difficult to solve, as it involves several OT sub-problems for which one of the marginals can change in the optimisation. An elegant approach leverages the fixed-point property $x=\sum_{i}\lambda_{i}T_{i}(x)$ which holds at the solution (under mild assumptions), where each $T_{i}$ is the OT map from $\nu$ to $\mu_{i}$ [1, 4]. Intuitively, this states that ‘each point in the support of the barycentre is the average of the corresponding points in the marginals’. This property has motivated an iterative fixed-point approach for barycentre computation, which involves iteratively updating a candidate barycentre $\nu$ by solving for the OT maps $T_{i}$ and constructing the next iterate as $\bar{T}\#\nu$ using the pushforward map $\bar{T}(x)=\sum_{i}\lambda_{i}T_{i}(x)$. Such methods have been shown to be highly successful in practice [4, 47, 55, 86].
@@ 2.2 Iterative Markovian Fitting
We now outline recent flow-based generative modelling approaches for solving for the SB problem, which will form the basis for our approach. The SB solution $\mathbb{P}^{SB}$ can be characterised as the unique path measure that is both Markov, and a mixture of bridges $\mathbb{P}^{SB}=\mathbb{P}_{0,T}^{SB}{\mathbb{Q}}_{\cdot|0,T}$, that has correct marginals $\mathbb{P}^{SB}_{0}=\mu_{0},\mathbb{P}^{SB}_{T}=\mu_{T}$ [60]. This property motivates the Iterative Markovian Fitting (IMF) procedure [77, 68], which solves for the SB solution by alternately projecting between Markovian processes and processes with the correct bridges. We refer to [77] for full details of the IMF procedure, but recall here the basic presentation. We recall the following definitions.
@@
A path measure $\Pi\in\mathcal{P}(\mathcal{C})$ is in the reciprocal class $\mathcal{R}({\mathbb{Q}})$ of $\mathbb{Q}$} if it is a mixture of bridges of $\mathbb{Q}$} conditional on their values at the endpoints, $\Pi=\Pi_{0,T}{\mathbb{Q}}_{\cdot|0,T}$. For a path measure $\mathbb{P}\in\mathcal{P}(\mathcal{C})$, the reciprocal projection is defined to be the mixture of bridges according to its induced coupling, $\operatorname{proj}_{\mathcal{R}({\mathbb{Q}})}(\mathbb{P})=\mathbb{P}_{0,T}{\mathbb{Q}}_{\cdot|0,T}$.
@@
It can be shown that, under mild conditions, these definitions coincide with the following minimisation problems over path measures [77],
@@
The IMF procedure requires learning the vector field corresponding the Markovian projections $\operatorname{proj}_{\mathcal{M}}(\Pi)$. This is done by training a neural network $v_{\theta}$ with a bridge-matching loss objective [67], for which the drift in (5) is the optimum,
@@
Traditionally, the standard way to solve the SB problem is via the Iterative Proportional Fitting (IPF) procedure [33, 51, 73], which minimises the KL divergence between subsequent iterations while alternating the endpoint measure that is fixed.
@@
This method was implemented using diffusion model-based approaches in [24], giving the Diffusion Schrödinger Bridge (DSB) method (see also [89], [12]). While allowing for arbitrary reference measures $\mathbb{Q}$}, this approach suffers from several shortfalls, such as only preserving both marginals at convergence (in comparison to IMF, which preserves both marginals at each iteration), as well as expensive trajectory caching, and ‘forgetting’ of the original reference measure [31]. As a result, IPF approaches have largely been superseded by IMF approaches which avoid these issues.
@@ 2.3 Tree-structured Schrödinger bridge
In this work, we consider the tree-structured Schrodinger Bridge problem [38]. In particular, we extend the IMF procedure to the dynamic Tree SB setting of [64].
@@
In order to explain the dynamic Tree SB problem, we first need to define stochastic processes on a tree. We will consider trees $\mathcal{T}=(\mathcal{V},\mathcal{E},\ell)$ with vertex set $\mathcal{V}$ and edge set $\mathcal{E}$, where $\ell:\mathcal{E}\rightarrow{\mathbb{R}}^{>0}$ is an edge-length function. One can extend the tree to a uniquely arcwise-connected metric space in the natural way; the arc connecting the two endpoints of an edge $e$ is identified with a line segment of corresponding length $T^{e}=\ell(e)$, and they are connected according to the graph structure. Via a slight abuse of notation we will also denote this metric space as $\mathcal{T}$. As such, we can define the space $C(\mathcal{T},{\mathbb{R}}^{d})$ of continuous paths from $\mathcal{T}$ into ${\mathbb{R}}^{d}$, and we let $\mathcal{P}(\mathcal{C}_{\mathcal{T}}):=\mathcal{P}(C(\mathcal{T},{\mathbb{R}}^{d}))$ denote the space of probability measures over such paths. We will present the methodology according to a directed tree $\mathcal{T}_{r}$ rooted at a vertex $r$, for which we can choose an ordered edge set $\mathcal{E}_{r}$ corresponding to a depth-first traversal (though we will often omit the dependence on $r$ in the notation). While this presentation may appear somewhat complex, the construction is quite natural; for an illustration of processes on a tree-structure, see Figures 1 and 4.
@@
As in the standard case, using the chain rule for KL divergence gives (under mild conditions) equivalence of the two problems; the dynamic SB solution can be expressed as a mixture of bridges over the static SB coupling, $\mathbb{P}^{SB}=\Pi_{\mathcal{V}}^{SB}{\mathbb{Q}}_{\cdot|\mathcal{V}}$. In this work, we aim to solve the $\mathrm{TreeSB}_{\text{dyn}}$ problem.
@@
This is precisely an entropy-regularised multi-marginal OT problem with a tree-structured quadratic cost function $c(x_{\mathcal{V}})=\sum_{(u,v)\in\mathcal{E}}\tfrac{1}{T^{(u,v)}}\lVert x_{u}-x_{v}\rVert_{2}^{2}$ and entropy-regularisation $\varepsilon=2\sigma^{2}$ (see Appendix A for an overview). Note that each weight is the reciprocal of the corresponding edge length.
Connection to Wasserstein barycentres
There are many ways to define entropy-regularised Wasserstein barycentres (see Appendix A for an overview). The work of [14] unifies several of these, combining $\varepsilon$-regularised OT costs with an entropic penalty term on the barycentre weighted by $\tau$. [64] show that the following doubly-regularised barycentre problem